OpenStudy (anonymous):

Simplify Each Product. 30.-6√15s^3 x 2√75

jhonyy9 (jhonyy9):

30-6s*sqrt(15s) * 2*5sqrt3 =30-6s*sqrt(15s) * 10sqrt3 = 30 - 60s*sqrt(45s)

OpenStudy (pfenn1):

@jhonyy9, I think the 30 is the problem number, it isn't part of the expression. \[-6\sqrt{15s^3} \times2\sqrt{75}=-6s \sqrt{15s} \times 10\sqrt{3}=-60s \sqrt{15s}\sqrt{3}=-180\sqrt{5}\]

jhonyy9 (jhonyy9):

yes you are right good luck bye and thank you for correction

OpenStudy (pfenn1):

you're welcome